Using your method of choice, what is the ordered pair?
x=7y+1
-3x+2y=-3
*Below is how far I've gotten before I got stuck:
-3(7y+1)+2y=-3
-21y-3+2y=-3
-19y-3=-3
-19y=0
Where do I go from here?
thanks!!!!!

Answer by ewatrrr(24785)


Hi
Using your method of choice, what is the ordered pair?
x=7y+1
-3x+2y=-3
*Below is how far I've gotten before I got stuck:
-3(7y+1)+2y=-3 |You have chosen substitution. Great choice!
-21y-3+2y=-3
-19y-3=-3
-19y = 0 |Good work thus far
y = 0 and x = 1 Ordered pair(1,0) the solution for this system
CHECKING our Answer***
-3*1 + 2*0 = -3